# PANTS: Practical Adversarial Network Traffic Samples against ML-powered Networking Classifiers

### v1

@@ 3.4 PANTS: AML for networking
To address the inherent challenges in applying AML to MNCs, we integrate a formal-method component to ensure the realizability and semantics-preservation of the produced adversarial samples. Our first insight is to use logical formulas to model the feature-engineering process that turns packet sequences into feature vectors, and sequence-level constraints for MNCs with no feature-engineering, effectively encoding input dependencies. Further, we encode threat-model capabilities, networking rules and semantics into flow-wide or packet-level constraints and ask an SMT (Satisfiability Modulo Theories) solver to find a packet sequence that satisfies them. If the answer is “UNSAT”, the adversarial features are not realizable under the given threat model, semantic, and networking constraints. Observe that SMT supports both forward and backward reasoning, hence can connect the requirements for adversariness with realizability and semantics. Critically, this makes PANTS extendible to any definition of semantics.
@@
As feature engineering modules, and sematics definition can be arbitrarily convoluted, a naive serial combination of AML with SMT will also fail because most of the AML-generated samples would not have a corresponding packet sequence that is realizable. Our second insight to address this challenge is that not all AML-generated features need to be honored for the sample to be adversarial. We design an optimization loop containing an AML component (e.g., PGD or ZOO), an SMT solver, and the target MNC, which iteratively checks whether a subset of the AML-produced adversarial features (or perturbations) can result in a realizable packet sequence until enough adversarial and realizable samples are collected. Instead of trying every possible subset of features, our third insight is to identify the most vulnerable adversarial features i.e., those that are both easy to tweak based on the threat model and important enough for the target ML model to affect its output. Critically, though, PANTS might opt out of constraints that make the generated samples adversarial but not constraints that make samples realizable or semantics preserving, which are always satisfied.
@@ 4.1 PANTS end-to-end view
We provide an end-to-end view of how PANTS generates adversarial packet sequences, shown in figure 4. The procedure includes 8 steps, starting from feeding in the original packet sequence to getting the adversarial and realizable packet sequence returned by PANTS.
@@ 4.2 SMT formulations
We leverage an SMT solver to help find adversarial packet sequences. We encode into formulas the feature dependencies, the threat model constraints, the networking and the semantics constraints. The SMT solver determines whether the resulting formulas are satisfiable.
@@
Equation (1), (2), (3) and (4) are forcing the adversarial packet sequences found to be consistent with the average (equation (1)), maximum (equation (2) and (3)) inter-arrival times and the average of packet length (equation (4)) in the adversarial features, maintaining the feature dependency. Equation (5) ensure that the packet sequence is complied with the threat model that it can only delay and append dummy payload to the forward packets. Equation (6) and  (7) formulate the network and semantics constraints.
@@ 4.3 Adversarial packet sequence identification
Feature importance determination.
@@
Considering all features will be too computationally expensive for the SMT solver and can lead to unsatisfiable combinations. Thankfully, not all features are equally important in keeping the packet sequences adversarial. Thus, we design a feature importance ranking algorithm and generate packet sequences that are only consistent with the most important features. Since AML methods are designed to perturb features to maximize the chance that they will be misclassified, our intuition is that the features that have been changed the most play a more important role in adversariness. Thus, after normalizing all the features into the same scale, the importance of each feature is determined by the difference in that feature value between the adversarial and the original one.
As an example, suppose there is a VPN flow p, and its extracted features are min and max of interarrival times, i.e., [$min_{iat}$, $max_{iat}$]. AML generates adversarial features to be $min^{\prime}_{iat}$ and $max^{\prime}_{iat}$ so that it is classified as Non-VPN. Unfortunately, SMT cannot find solutions consistent with these two features. Therefore, we want to select the most important feature and find the packet sequences that only comply with that one. Assume that $min^{\prime}_{iat}$ is more important than $max^{\prime}_{iat}$. Then, the SMT solver finds solutions only for $min^{\prime}_{iat}$. After that, we construct the packet sequences using these solutions and recalculate the maximum interarrival times of the given packet sequences $max^{\prime\prime}_{iat}$. Since $min^{\prime}_{iat}$ is consistent with AML’s decision and $min^{\prime}_{iat}$ is the most important feature, we expect that the new features [$min^{\prime}_{iat}$, $max^{\prime\prime}_{iat}$] will still be adversarial.
Find adversarial and realizable packet sequences using iterative testing. After determining the important features, the next questions are: (i) how many important features are necessary so that the conforming packet sequences are adversarial; and (ii) how to ensure that the SMT solver can find packet sequences consistent with these selected important features? To address them, we employ iterative testing, which searches for successful adversarial packet sequence generation, starting with the most important features. Starting with an empty feature list, we iteratively test whether adding the formulas of the next most important (not already explored) feature $f_{i}$ results in a conforming packet sequence output from the SMT solver. If it does, $f_{i}$ is appended to the selected features list. Otherwise, the feature is not selected. By adding more and more features, the found packet sequences are more likely to be adversarial. In practice, it is possible that after appending $f_{i}$, the solution provided by the SMT solver is already adversarial. However, in order to cover a wider spectrum of adversarial features and enable robustification of the model, the algorithm collects this solution and continues appending $f_{i+1}$ for another potential solution until $i$ reaches $k$. $k$ is a parameter defined by the users, representing the $k$ most important features that need to be included in the iterative testing. Setting a small value forces the algorithm to include only a few features, which can hurt the adversariness. Setting it too large results in including too many non-important features, making the generation speed slow without much adversariness improvement. The algorithm details are depicted in alg 1.
@@
original_sample $\leftarrow$ the original and benign packet sequence
sorted_feature_list $\leftarrow$ ($f_{1}$, $f_{2}$, …, $f_{n}$) $\triangleright$ Sorted by importance descending order
selected_features $\leftarrow$ ()
solutions_set $\leftarrow$ ()
function SMT_solver(selected_features, original_sample)
    if  SMT can find solutions consistent with selected_features then
       return solutions
    else
       return null
    end if
end function
for $i$ in 1:k do $\triangleright$ $k\leq n$
    selected_features.append($f_{i}$)
    solution $\leftarrow$ SMT_solver(selected_features, original_sample)
    if solution is null then
       selected_features.remove($f_{i}$)
    else
       solutions_set.append(solution)
    end if
end for
Check adversariness for all the solutions in solutions_set
@@ 4.4 PANTS’ SMT optimizations
The SMT solver’s efficiency in finding adversarial packet sequences is a critical factor for PANTS’ usability. Unfortunately, the SMT solver is often unable to find adversarial packet sequences within a reasonable time frame. The problem is that SMT solvers have scalability limitations and the given formulas can result in very high computational complexity. The complexity depends on many factors, such as the length of the packet sequence, the formulation of capabilities, the number of features to be solved, and the difficulty of solving each feature. We have already discussed restricting the number of used features and prioritizing the evaluation of the most critical features in §4.3. In this section, we focus on optimizing the packet sequence length and the formulation of capabilities.
@@
We first set a time budget for the SMT solver. If the SMT solver cannot find packet sequences within a timeout window, we decrease the computation complexity by dividing a flow into several chunks evenly. The SMT solver is used to find the packet sub-sequences for each chunk. Since each chunk has fewer packets than the whole flow, it is more likely that the SMT solver will be able to find solutions within the timeout limits. After splitting a flow into multiple chunks, some features are difficult to be formulated strictly. For example, when working on a complete flow, it is easy to formulate the maximum packet length feature. After splitting the flow into multiple chunks, the maximum packet length can appear in any chunk. Strictly formulating this feature would require a complicated algorithm, which is in conflict with our motivation to decrease computation complexity and is unnecessary for our use case. To solve this problem, we consider three different types of features. For the first type of features, satisfying their formulas for each sub-sequence will likely result in satisfied formulas for the whole sequence. For example, such features are {max, mean, median, min, std} of {interarrival times, packet length} and number of {packets, bytes} per second. For these features, SMT solves their formulas per chunk. For the second feature type, the accumulation of each chunk’s feature values is equal to that of the whole flow. One typical example is the total duration. For these features, the SMT solver is finding sub-sequences in each chunk consistent with $\frac{feature}{\#chunks}$. For other features, such as the number of unique packet lengths, we are not aiming to be consistent with the whole-flow formulas when finding the sub-sequences for each chunk.
Finally, parallel computing can further speed up generation time, since the SMT-solving process for each flow/chunk is designed to be independent.
Figure 5: The average elapsed time for PANTS generating a single sample under different ML models, threat models and applications. In most of the cases, generating one sample takes less than 1 second.
@@
Figure 5 shows the sample generation speed after the discussed optimizations. PANTS achieves reasonable efficiency by taking less than 1s on generating a single sample in most of the cases. Generally, the generation speed doesn’t show notable difference between threat models, ML models and applications. However, for MLP and RF tested for the in-path threat model, the generation speed varies across different applications. The in-path threat model indicates a limited action space, making SMT harder to find solutions especially for those processing pipelines using feature engineering.
4.5 Adversarial Training with PANTS
Adversarial training is often used to enhance the robustness of ML models. During authentic adversarial training, an AML method (e.g., PGD) replaces each benign sample with an adversarial one for the insufficiently trained version of the model. While useful, this method can hurt the accuracy of an ML model [17, 31] and only applies to NN-based models. To avoid these shortcomings, we adopt an alternative way of adversarial training, in which we augment the training dataset with adversarial samples generated for the sufficiently-trained model to fine-tune it. This methodology is model agnostic, meaning that it can work with any processing pipeline. We also find in §5.2 that this training methodology improves robustness without hurting the model’s accuracy.
@@ 5 Evaluation
In this section, we evaluate the efficacy of PANTS in two ways. First, we evaluate PANTS’ capability in generating adversarial realizable, and semantic-preserving flows and compare it with two baselines, Amoeba [26] and BAP [29]. Second, we evaluate the effectiveness of PANTS-generated samples in improving the robustness of the targeted MNCs. We begin by introducing the evaluation setup, and then we discuss some key findings. In summary, PANTS is 70% more likely to find adversarial samples compared to Amoeba in median and 2x more likely compared to BAP. Using PANTS for robustification can have 142% more robustness improvement then using Amoeba without sacrificing the model accuracy.
@@ 5.1 Setup
Applications.  We select three typical classification tasks to evaluate PANTS. They have diverse objectives, flow statistical patterns, and classification granularity, providing a representative sample of networking applications.
@@
VPN traffic detection is a challenging task for encrypted traffic classification. A well-functioning classifier is trained to determine whether the traffic is VPN. ISCXVPN2016[11] is one of the commonly-used labeled datasets, and includes both VPN and non-VPN traffic. The dataset contains more than $20$ million packets grouped into $8,577$ bi-directional flows. Of those, $71$% are non-VPN, and the rest are VPN flows. This is an unbalanced dataset, hence, we include both accuracy and F1 score for more comprehensive evaluation of the model performance.
@@
Application identification (APP) is useful for characterizing mobile traffic from different applications and is crucial for network management tasks such as anomaly detection and resource allocation. UTMobileNetTraffic2021 [18] is a commonly-used dataset that includes mobile traffic from 18 popular applications (e.g., Dropbox, Google Drive, Facebook) under different user activities (e.g., Dropbox-download, Google Drive-upload, Facebook-scroll). It has multiple subsets representing different controlled user activities, such as scrolling the Facebook news feed. We choose one of these subsets that collects application traffic without placing any constraints on user activity. Similar to VPN detection, classification is applied to the bi-directional flows. The selected dataset contains 7,134 bi-directional flows.
@@
Quality of Experience (QoE) inference arises as an important task for all stakeholders, such as ISPs and service providers, due to the rise in the use of Video Conferencing Applications (VCAs). By training various models, multiple QoE metrics can be inferred, e.g., resolution and video bitrate directly from packet traces. The state-of-the-art framework is trained and tested on a dataset called VCAML in a per-second granularity[33]. VCAML contains a large number of single-directional video conferencing flows from Google Meet, Microsoft Teams and Cisco Webex in different resolutions, bitrate, etc. We pick resolution prediction as the task to test our framework. The total number of samples is 37,274, which includes 11 different resolutions, such as 720p, 360p.
The features extracted by the feature engineering module for different applications are listed in the Appendix in  §A. All the datasets are applied with basic data cleaning. Any flow with only one packet is removed as calculating interarrival times requires at least two packets.
@@
ML models.  For each application, we train four commonly used ML classifiers independently. They are Multilayer Perceptron (MLP), Random Forest (RF), Transformer (TF) and Convolutional Neural-Network (CNN). They exemplify different ML approaches, including (deep) neural network-based and tree-based techniques. Meanwhile, they also represent different processing pipelines. MLP and RF use a non-differentiable and non-invertible feature engineering module, which extracts statistical features that are then used as input to the models to predict the labels. TF and CNN are end-to-end differentiable: packet sequences are directly encoded as a sequence of interarrival times, packet direction and packet length for classification. The maximum sequence length is set to 400. The details of the hyperparameters used for each model are explained in the Appendix in §B. Table 1 shows the performance of these 4 ML classifiers. For each dataset, the training set ($D_{train}$) and the testing set ($D_{test}$) are following a 80%:20% random split. ML models are trained on $D_{train}$, and the performance is evaluated on $D_{test}$.
Application Model Accuracy F1 Score
APP MLP 0.7947 0.7966
RF 0.8206 0.8202
TF 0.7666 0.7645
CNN 0.7926 0.7923
VPN MLP 0.8902 0.9437
RF 0.9400 0.9574
TF 0.9295 0.9287
CNN 0.9219 0.9214
QOE MLP 0.8070 0.8247
RF 0.8221 0.8330
TF 0.7819 0.7831
CNN 0.8203 0.8159
Table 1: Performance of vanilla ML classifiers for each application. For multi-class classification, we calculate the F1 score per class in a one-vs-rest manner. We apply weighted averaging afterwards to have a single score for easier comparison. Our trained models demonstrate reasonably good accuracy and F1 score for each application.
Attacker’s capability.  To show the flexibility of PANTS, we evaluate two realistic threat models. The two models correspond to attackers located at the end host and in-path, respectively. The attackers at the end host are stronger as they can delay packets, inject packets, and append dummy payload to the existing packets. To preserve the flow’s semantics, we consider that the end-host attacker can only delay the flow up to 20% of the original flow’s duration, inject at most 20 packets, and append dummy payload to at most 20% of the packets. Of course any flow-wide or packet-level constraint would work. The in-path attackers are weaker as they can only delay packets. An in-path attacker can only delay at most 20% of the original flow’s duration. Note that our assumption considers a single attacker at a time, thereby restricting the manipulation to one direction of flow. Adding more threat models to PANTS consists of adding the relevant constraints.
Baselines.  Since there is no related framework that aims to help the operator assess and enhance the robustness of MNCs, we compare them against the State-Of-The-Art attack generation methods, Amoeba and BAP. Amoeba [26] is an RL-based proposal to attack ML-based censorship. Whilst designed for a different goal, Amoeba also generates adversarial samples against applications that are end-to-end differentiable and that include a non-differential or non-invertible module. Amoeba supports delaying packets, appending dummy payload to the existing packets, and splitting packets into multiple ones. For a fair comparison, we extend Amoeba’s capability to comply with the two evaluated threat models. We also train Amoeba’s agent to respect the flow’s semantics and enforce a budget during inference. We explain the details of our extension in the Appendix in §C.
BAP is to train a generator to create perturbations for each packet. Since BAP [29] does not support applications that include non-differential and non-invertible components, we only consider it for TF and CNN but not MLP and RF. The original implementation of BAP can only support perturbing every packet, so we extend BAP by adding a mask during training and inference to ensure reliability (e.g., only act on forward packets). Although BAP supports per-packet constraints, it does not support flow-wide constraints such as those needed for semantics preservation.
Testbed.  Appendix §D details the used CPU/GPU machines.
5.2 Key Findings
Finding 1: PANTS is 70% more likely to find adversarial samples compared to Amoeba in median and 2x more likely compared to BAP.
@@ 5.2 Key Findings
Figure 6: The attack success rate (ASR) of PANTS, Amoeba and BAP for various ML models, applications and threat models. PANTS has a much higher ASR compared to Amoeba and BAP, demonstrating its ability to effectively generate adversarial samples, which are then used for debugging, fine-tuning, and robustness assessment.
We first compare PANTS with Amoeba and BAP in their ability to generate adversarial flows for various ML models, applications, and threat models. We run experiments five times and report the Attack Success Rate (ASR), i.e., the ratio of realizable, semantic-preserving and adversarial samples / total original samples for each technique [^2] While PANTS generates multiple adversarial samples per original, for the ASR calculation each original sample can contribute at most one adversarial.. Figure 6 shows the results of ASR. PANTS achieves 35.31% ASR in median while only 19.57% for Amoeba and 10.31% for BAP. Among these three methods, BAP performs the worst in part due to its inability to respect semantics i.e., flow-level constraints.
Compared to Amoeba, PANTS is 51% more likely to find adversarial samples for the end-host threat model in median and 2X more likely for in-path attacks. We observe that Amoeba struggles to find adversarial samples when for an attacker that is in-path. Concretely, for the QoE inference, the ASR for Amoeba is around 0.3% for in-path threat model while PANTS is 37%. PANTS is particularly more effective with the weak threat model compared to Amoeba, showing its more comprehensive nature.

### v3

@@ 3.4 PANTS: Adversarial ML for networking
To address the inherent challenges in applying AML to MNCs, we integrate a formal-method component to ensure the realizability and semantics-preservation of the produced adversarial samples. Our first insight is to use logical formulas to model the feature-engineering process that turns packet sequences into feature vectors and sequence-level constraints for MNCs with no feature-engineering, effectively encoding input dependencies. Further, we encode threat-model capabilities, networking rules and semantics into flow-wide or packet-level constraints and ask an SMT (Satisfiability Modulo Theories) solver to find a packet sequence that satisfies them. If the answer is “UNSAT”, the adversarial features are not realizable under the given threat model, semantic, and networking constraints. Observe that SMT supports both forward and backward reasoning, hence can connect the requirements for adversariness with realizability and semantics. Critically, this makes PANTS extensible to any definition of semantics.
@@
As feature engineering modules and semantics definition can be arbitrarily convoluted, a naive serial combination of AML with SMT will also fail because most of the AML-generated samples would not have a corresponding packet sequence that is realizable. Our second insight to address this challenge is that not all AML-generated features need to be honored for the sample to be adversarial. We design an optimization loop containing an AML component (e.g., PGD or ZOO), an SMT solver, and the target MNC, which iteratively checks whether a subset of the AML-produced adversarial features (or perturbations) can result in a realizable packet sequence until enough adversarial and realizable samples are collected. Instead of trying every possible subset of features, our third insight is to identify the most vulnerable adversarial features i.e., those that are both easy to tweak based on the threat model and important enough for the target ML model to affect its output. Critically, though, PANTS might opt out of constraints that make the generated samples adversarial but not constraints that make samples realizable or semantics preserving, which are always satisfied.
@@ 4.1 PANTS end-to-end view
We provide an end-to-end view of how PANTS generates adversarial packet sequences, shown in Fig. 4. The procedure includes 8 steps, starting from feeding in the original packet sequence to getting the adversarial and realizable packet sequence returned by PANTS.
@@ 4.2 SMT formulations
PANTS incorporates an SMT solver for generating adversarial and realizable packet sequences. Concretely, the SMT solver encapsulates constraints for feature dependencies, the threat model, networking rules, and semantics, determining their satisfiability, i.e., whether a sample satisfying these constraints exists.
@@
For this example, we thereby define four variables $\Delta t_{1}^{\prime}$, $\Delta t_{3}^{\prime}$, $l_{1}^{\prime}$ and $l_{3}^{\prime}$. They are representing the delay and the appended payload for $p_{1}^{+}$ and $p_{3}^{+}$. The perturbed flow p’ can be formulated as p’=$=\{\bar{p}_{1}^{+},p_{2}^{-},\bar{p}_{3}^{+}\}=\{(\Delta t_{1}+\Delta t_{1}^{\prime},l_{1}+l_{1}^{\prime})^{+},(\Delta t_{2},l_{2})^{-},(\Delta t_{3}+\Delta t_{3}^{\prime},l_{3}+l_{3}^{\prime})^{+}\}$. Then, the formula for the SMT solver is
@@
Equation (1), (2), (3) and (4) force the generated packet sequences to be adversarial by being consistent with the average (equation (1)), maximum (equation (2),(3)) inter-arrival times and the average of packet length (equation (4)) that AML specification. Equation (5) ensures that the packet sequence complies with the threat model, which includes delaying and appending dummy payload to the forward packets. Equation (6) and  (7) formulate the network and semantics constraints.
@@ 4.3 Adversarial packet sequence identification
Feature importance determination(line 1 - line 5 in Alg. 1).
@@
Adhering to all features in the adversarial feature vector that the AML specified could be too computationally expensive for the SMT solver or even infeasible. Critically, though, not all features are equally important in keeping the packet sequences adversarial. For instance, given an adversarial vector of two concrete values generated by the AML component, there is often a packet sequence that adheres to only one of them while still being adversarial. Building on this insight, we design a feature importance ranking algorithm that helps PANTS identify feature subsets that play a more important role in the sample’s adversariness, thereby streamlining and accelerating the search process. Since AML methods are designed to perturb features to maximize the chance that they will be misclassified, the features that have been changed the most play a more important role in adversariness. Thus, after normalizing all the features, the importance of each feature is determined by the difference between the adversarial and the original value of that feature. Using this importance ranking, we can prioritize adversarial-feature constraints, which are crucial to keeping the flow adversarial.
Find adversarial and realizable packet sequences using iterative testing (line 19 - line 28 in Alg. 1).
After quantifying feature importance, PANTS leverages this information to maximize the likelihood of quickly identifying adversarial and realizable samples. Rather than starting with the full set of adversarial-features constraints or testing random permutations—both of which would significantly increase computational overhead in the SMT solver—PANTS employs the following algorithm to optimize the process.
(a) End-host attacker
(b) In-path attacker
Figure 5: The impact of $k$ in ASR (i.e., in PANTS ability to generate examples) against MLP and RF. The ASR increases with $k$, but levels off once $k$ reaches a certain threshold.
Starting with an empty feature list, we iteratively test whether adding the formulas of the next most important (not already explored) feature $f_{i}$ results in a conforming packet sequence output from the SMT solver. If it does, $f_{i}$ is appended to the selected features list. Otherwise, the feature is not selected. By adding more and more features, the found packet sequences are more likely to be adversarial. Even after the SMT solver finds a packet sequence that is adversarial, the algorithm continues appending more features until it appends $k$ features, to cover a wider spectrum of adversarial examples. The exact value of $k$ can be set and easily fine-tuned by the user. Concretely, PANTS finds more diverse adversarial examples as one increases $k$ (up to a point), as we observe in Fig. 5, but will also take longer to compute them. In practice, one only needs to avoid setting $k$ too low, making half the number of features a reasonable default.
Thanks to these two key logics, PANTS is able to identify adversarial packet sequences. The algorithm details are depicted in Alg. 1.
@@
[…]
4:solutions_set
5:function Feature_Rank(original_sample, adv_features)
6:    original_feature = feature_engineer(original_sample)
7:    sorted_feature_list = sort(abs(adv_features - original_features))
[…]
9:end function
10:
11:function SMT_Solver(selected_features, original_sample)
12:    if  SMT can find solutions consistent with selected_features then
13:       return solutions
14:    else
15:       return null
16:    end if
17:end function
18:
19:sorted_feature_list $\leftarrow$ ()
20:selected_features $\leftarrow$ ()
21:solutions_set $\leftarrow$ ()
22:sorted_feature_list $\leftarrow$ Feature_Rank(original_sample, adv_features)
23:for $i$ in 1:k do
24:    selected_features.append(sorted_feature_list[$i$])
25:    solution $\leftarrow$ SMT_Solver(selected_features, original_sample)
26:    if solution is null then
27:       selected_features.remove(sorted_feature_list[$i$])
28:    else
29:       solutions_set.append(solution)
30:       Check adversariness for this solution
31:    end if
32:end for
@@ 4.4 PANTS’ SMT optimizations
While preferentially enforcing subsets of adversarial features improves scalability, there are scenarios where the SMT solver’s limitations become a more difficult bottleneck. For example, there will be cases in which finding an adversarial example requires adhering to all feature constraints or cases where the most significant feature constraint is so complex that the solver struggles. The complexity depends on many factors, such as the length of the packet sequence, the formulation of capabilities, the number of features to be solved, and the difficulty of solving each feature. In this section, we focus on optimizing the packet sequence length and the formulation of capabilities.
@@
We first set a time budget for the SMT solver. If the SMT solver cannot find packet sequences within a timeout window, we decrease the computation complexity by dividing a flow into several chunks evenly. The SMT solver is used to find the packet sub-sequences for each chunk. Since each chunk has fewer packets than the whole flow, it is more likely that the SMT solver will be able to find solutions within the timeout limits. After splitting a flow into multiple chunks, some features are difficult to formulate strictly. For example, when working on a complete flow, it is easy to formulate the maximum packet length feature. After splitting the flow into multiple chunks, the maximum packet length can appear in any chunk. Strictly formulating this feature would require a complicated algorithm, which is in conflict with our motivation to decrease computation complexity and is unnecessary for our use case. To solve this problem, we consider three different types of features. For the first type of features, satisfying their formulas for each sub-sequence will likely result in satisfied formulas for the whole sequence. For example, such features are {max, mean, median, min, std} of {interarrival times, packet length} and number of {packets, bytes} per second. For these features, SMT solves their formulas per chunk. For the second feature type, the accumulation of each chunk’s feature values is equal to that of the whole flow. One typical example is the total duration. For these features, the SMT solver is finding sub-sequences in each chunk consistent with $\frac{feature}{\#chunks}$. For other features which exhibit convoluted correlations between the sub-sequences and the whole sequence, such as the number of unique packet lengths, we are not aiming to be consistent with the whole-flow formulas when finding the sub-sequences for each chunk. The number of chunks has minimal impact on ASR. Indeed, we found that even doubling their number (from 10 to 20 chunks) changes ASR by 3%. That is because once the solver is able to find sub-sequences for each chunk, continuing to split into smaller chunks does not improve ASR.
Finally, PANTS’ use of the solver is highly parallelizable, facilitating further speed ups. Indeed, the solver for each original input, chunk, and even set of adversarial features of a given instance can run independently.
@@
Fig. 12 shows the sample generation speed after the discussed optimizations. PANTS generates ~1.7 samples/sec, which is efficient, considering MNCs are rarely re-trained.
4.5 Iterative Adversarial Training with PANTS
Adversarial training is often used to enhance the robustness of ML models. During authentic adversarial training, an AML method (e.g., PGD) replaces each benign sample with an adversarial one for the insufficiently trained version of the model. While useful, this method can hurt the accuracy of an ML model [25, 46] and only applies to NN-based models. To avoid these shortcomings, we design an alternative way of adversarial training, in which we augment the training dataset with adversarial samples generated for the sufficiently trained model to fine-tune it. This methodology is model-agnostic, meaning that it can work with any processing pipeline. We also find in §5.2 that this training methodology improves robustness without hurting the model’s accuracy.
For robustification, PANTS generates multiple adversarial samples for each benign counterpart in total to cover diverse adversarial samples. Using PANTS, we generate around 1.16-180 samples for each benign sample during robustification based on different threat models, applications and ML models. Along with adversarial samples, we include non-adversarial ones generated by PANTS. This is beneficial for training because the non-adversarial samples are closer to the decision boundary compared to their benign counterpart, despite being correctly classified. Including those samples can help form a more refined decision boundary.
@@ 5 Evaluation
In this section, we evaluate PANTS in two ways. First, we evaluate PANTS’ capability in generating adversarial realizable and semantic-preserving flows and compare it with two baselines, Amoeba [35] and BAP [41]. We find that PANTS is 70% more likely to find such flows compared to Amoeba in median and 2x more likely compared to BAP. Second, we evaluate the effectiveness of PANTS-generated samples in improving the robustness of the targeted MNCs. We find that PANTS is 142% more effective in improving robustness compared to Amoeba without sacrificing the model accuracy. Interestingly, we empirically find PANTS enhances the model’s robustness even against attackers that have different (and even strictly more) capabilities than what was assumed during sample generation. We begin by introducing the evaluation setup before discussing some key findings.
@@ 5.1 Setup
Applications & Datasets.  We evaluate PANTS in three MNCs with diverse objectives, flow statistical patterns, and classification granularity, providing a representative sample of networking applications. We evaluate several implementations of each MNC, including the exact implementation described in the paper from which each dataset is sourced.
@@
VPN: VPN traffic detection aims at identifying flows that are VPN among encrypted traffic flows[18, 43, 21, 38]. ISCXVPN2016 [18] is one of the commonly used labeled datasets and includes both VPN and non-VPN traffic. The dataset contains more than $20$ million packets grouped into $8,577$ bi-directional flows. Of those, $71$% are non-VPN, and the rest are VPN flows.
@@
APP: Application identification aims at predicting the application that given traffic serves and is useful for network management tasks such as anomaly detection and resource allocation [39, 26, 36, 52, 54]. UTMobileNetTraffic2021 [26] is a commonly used dataset that includes mobile traffic from 18 popular applications (e.g., Dropbox, Google Drive, Facebook). The selected dataset contains 7,134 bi-directional flows.
@@
QoE: Quality of Experience inference aims at inferring QoE metrics like resolution and video bitrate from packet traces[48, 13, 42, 59, 27]. QoE inference is useful for stakeholders, such as ISPs and service providers, due to the rise in the use of Video Conferencing Applications (VCAs). VCAML [48] is the dataset that contains a large number of single-directional video conferencing flows in a per-second granularity from Google Meet, Microsoft Teams, and Cisco Webex. The total number of samples is 37,274, which includes 11 different resolutions, such as 720p, 360p.
The features extracted by the feature engineering module for different applications are listed in the Appendix in §B.
@@
Application implementations.  For each application, we train four commonly used ML classifiers independently. They are Multilayer Perceptron (MLP), Random Forest (RF), Transformer (TF) and Convolutional Neural-Network (CNN) to cover multiple ML approaches, and processing pipelines (i.e., end-to-end differentiable or not). The details of the hyperparameters used for each model are explained in the Appendix in §C. Table 1 shows the performance of these three MNCs. For each dataset, the training set ($D_{train}$) and the testing set ($D_{test}$) are following a 80%:20% random split. ML models are trained on $D_{train}$, and the performance is evaluated on $D_{test}$.
Application
Accuracy
Accuracy
(In-path/End-host robustified)
APP [0.7666, 0.8206] [0.7400, 0.8136]
VPN [0.8902, 0.9400] [0.8596, 0.9375]
QOE [0.7819, 0.8221] [0.7898, 0.8173]
Table 1: Performance of vanilla ML classifiers for each application. Our trained models demonstrate reasonably good accuracy for each application. After robustification via PANTS, the robustified ones’ accuracy doesn’t drop.
@@
Figure 6: The attack success rate (ASR) of PANTS, Amoeba and BAP for various ML models, applications and threat models. PANTS has a much higher ASR compared to Amoeba and BAP, demonstrating its ability to effectively generate adversarial samples, which can be used for debugging, fine-tuning, and robustness assessment.
Attacker’s capability.  We evaluate two threat models that correspond to attackers located at the end host and in-path. The attackers at the end host (i.e., server or client) can delay packets, inject packets, and append dummy payload to the existing packets of one direction. To preserve the flow’s semantics, we consider that the end-host attacker can only delay the flow up to 20% of the original flow’s duration, inject at most 20 packets, and append dummy payload to at most 20% of the packets. The in-path attackers can only delay packets of one direction at most 20% of the original flow’s duration. We selected these two models because they represent realistic and rational attackers, but thanks to PANTS flexibility one can easily add more threat models by adding the relevant constraints as we do in Fig. 11.
Baselines.  Since there is no related framework that aims to help the operator assess and enhance the robustness of MNCs, we compare PANTS against the SOTA adversarial generation methods, Amoeba and BAP and a synthetic data generator, NetShare.
Amoeba [35] is an RL-based proposal to attack ML-based censorship. Although designed for a different goal, Amoeba also generates adversarial samples against applications that are end-to-end differentiable and that include a non-differential or non-invertible module. Amoeba supports packet delay, dummy payload appending, and packet split into multiple ones. For a fair comparison, we extend Amoeba’s capability to comply with the two evaluated threat models. We also train Amoeba’s agent to respect the flow’s semantics. We explain the details of our extension in the Appendix in §D.
BAP [41] is to train a generator to create perturbations for each packet. Since BAP does not support applications that include non-differential and non-invertible components, we only consider it for TF and CNN (i.e., not with MLP and RF). The original implementation of BAP can only support perturbing every packet, so we extend BAP by adding a mask during training and inference to ensure reliability (e.g., only act on forward packets). Although BAP supports per-packet constraints, it does not support flow-wide constraints such as those needed for semantics preservation.
NetShare [62] is a GAN-based synthetic data generator specialized in producing high-fidelity network packet traces. After NetShare is trained on raw packet sequences, it can generate synthetic ones that closely match the distribution of the original data. NetShare can successfully match the convoluted data distribution from backbone traces such as CAIDA [4] and data center traces such as UNI1 [11].
Success Metric.  To quantify the robustness of MNCs, we report the Attack Success Rate (ASR), defined as the ratio of realizable, semantic-preserving adversarial samples to the total original samples for each technique, following standard practice [12]. ASR provides a meaningful measure of robustness, as it directly evaluates a system’s resilience under stress and adversarial conditions. We also report accuracy for completeness and to assess performance under typical conditions. Accuracy refers to the proportion of successful classifications on the original test set (not adversarial), which is drawn from the same data source (hence the same distribution) as the training set. [^2] High ASR is orthogonal to high accuracy. MNCs can both have high accuracy and ASR, meaning that they perform well under standard conditions (high accuracy), but are susceptible to attacks (high ASR).
Testbed.  Appendix §E details the used CPU/GPU machines.
5.2 Key Findings
  1 PANTS is 70% more likely to find adversarial samples compared to Amoeba in median and 2x more likely compared to BAP.
We compare PANTS with Amoeba and BAP in their ability to generate adversarial flows under two threat models against three MNCs implemented with four models and report the ASR. As we observe in Fig. 6, PANTS clearly outperforms baselines in all cases. More specifically, PANTS achieves a 35.31% ASR (median across all cases), whereas Amoeba and BAP achieve 19.57% and 10.31%, respectively. BAP’s inefficiency stems from its inability to incorporate semantic constraints, e.g., flow-level constraints. Amoeba outperforms BAP (despite being a black-box approach), but PANTS is still 51% more likely than Amoeba to identify adversarial samples under the end-host threat model (median) and 2x more likely under the in-path threat model. This highlights that while Amoeba is effective at uncovering adversarial examples in scenarios with abundant opportunities (i.e., for stronger attackers), it struggles when such examples are less apparent.